OpenStudy (anonymous):

7x what and 8x what = the same thing

OpenStudy (anonymous):

0

OpenStudy (anonymous):

I guess it must be 0

OpenStudy (dumbcow):

7x = 8x subtract 7x on both sides 0=x

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i cant word it right

OpenStudy (dumbcow):

reciprocal ?? 7*1/7 = 1 8*1/8 = 1

OpenStudy (anonymous):

EX.3x5=15 and 5x3=15 what is this but with 7 and 8

OpenStudy (dumbcow):

7*8=56

OpenStudy (anonymous):

in that case it is 8 and 7 respectively

OpenStudy (anonymous):

which makes 7 *8 =56 and 8*7=56

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

4x8=

OpenStudy (anonymous):

um i need ajaydkulkarni

OpenStudy (anonymous):

for a question

OpenStudy (dumbcow):

4*8=32

OpenStudy (anonymous):

thax

OpenStudy (anonymous):

3x6=

OpenStudy (dumbcow):

you should know these :) multiplication facts also you could use a calculator

OpenStudy (anonymous):

sorry in only in 4 grade

OpenStudy (dumbcow):

oh haha sorry but still... ok 6*3 = 6+6+6